OpenStudy (anonymous):

Find the length of Acr AB: if the angle measure is 65 degrees and the diameter is 14in

OpenStudy (anonymous):

\[\frac{AB}{Perimeter}=\frac{65^o}{360^o}\rightarrow AB=\pi{14} \times \frac{65}{360} \approx7.94inches\]